OpenStudy (kainui):

Just looking for any and all cute closed/alternate forms for arithmetic functions.

OpenStudy (kainui):

Here's one I found on my own: von Mangoldt function: \[\Lambda(n) = \ln[ rad(n) \delta(\omega(n)) ] \] Functions used for the curious: rad(n) Basically makes all the exponents on the prime factorization =1. https://en.wikipedia.org/wiki/Radical_of_an_integer \(\delta(n) = \lfloor \tfrac{1}{n} \rfloor \) (In other words, 1 if n=1, 0 otherwise for positive integers) \(\omega (n)\) number of distinct prime factors http://mathworld.wolfram.com/DistinctPrimeFactors.html

OpenStudy (amistre64):

Does this count?

OpenStudy (amistre64):

forbidden script? lol ... well that was unexpected

ganeshie8 (ganeshie8):

.

OpenStudy (amistre64):

copy to notebook and save as html file .... <html> <body> <input type=button value=" Start " onCLick="potion()"> <div style="position:absolute; top:1; left:1" id="holder"></div> <div style="position:absolute; top:1; left:1" id="leftside"></div> </body> <script language=javascript> var g1, g2, g3, g4 var outP="" var l1= Math.PI*(-1) var l2= 0 var l3= 0 var l4= 0 var x1,y1,vx2,vy2, flag=0 function potion() { g1 = setInterval("topRight()", 42) } function topRight() { if(l1<=Math.PI/5) { vx2 = 250+(75*Math.cos(l1)) vy2 = 100+(75*Math.sin(l1)) outP=outP+"<di"+"v style='position:absolute; top:"+vy2+"; left:"+vx2+"'><fon"+"t color='red'><fon"+"t color='red'>*</f"+"ont></f"+"ont></div>" x1= vx2 y1= vy2 flag=1 } else { if (flag==1){g2=setInterval("topLeft()",42)} flag=0 vx2 = x1-l2 vy2 = y1+l2 outP=outP+"<di"+"v style='position:absolute; top:"+vy2+"; left:"+vx2+"'><fon"+"t color='red'>*</f"+"ont></div>" l2 = l2+1.5 if(l2>136) { clearInterval(g1) } } l1=l1+Math.PI/90 document.getElementById("holder").innerHTML=outP } var outL="" var x2,y2,vx1,vy1 function topLeft() { if(l3>=-6*Math.PI/5) { vx1 = 100+(75*Math.cos(l3)) vy1 = 100+(75*Math.sin(l3)) outL=outL+"<di"+"v style='position:absolute; top:"+vy1+"; left:"+vx1+"'><fon"+"t color='red'><fon"+"t color='red'>*</f"+"ont></f"+"ont></div>" x2= vx1 y2= vy1 } else { vx1 = x2+l4 vy1 = y2+l4 outL=outL+"<di"+"v style='position:absolute; top:"+vy1+"; left:"+vx1+"'><fon"+"t color='red'>*</f"+"ont></div>" l4 = l4+1.5 if(l4>136) {g3=setInterval("initials()",1500) clearInterval(g2) } } l3=l3-Math.PI/90 document.getElementById("leftside").innerHTML=outL } </script> </html>

OpenStudy (kainui):

Hahahhaa @amistre64 \[\huge \color{red} \heartsuit\]

OpenStudy (kainui):

Hahahhaa @amistre64 \[\huge \color{red} \heartsuit\]

OpenStudy (amistre64):

:)

OpenStudy (kainui):

Here's another one I just found, it's not particularly crazy or serious haha: \[rad( n!) = n \#\] \(n \#\) is the primorial, it's like a factorial only it is the product of all primes less than or equal to n.

OpenStudy (kainui):

Just discovered a new thing pretty cute: \[rad^2(n) = \sum_{d|n} \mu^2(d)\varphi(d) \sigma(d)\]

OpenStudy (kainui):

Just defined a new arithmetic function for fun: for n defined this way: \[n = \prod_{primes} p^{k}\] My new multiplicative function is: \[\chi(n) = \prod_{primes} p^{kp}\] and taking the arithmetic derivative of this function seems to give: \[\chi'(n) = \Omega(n)\chi(n)\] Also it appears that \(\chi(n)\) has an inverse, \[\chi^{-1}(n) = \prod_{primes} p^{k/p}\] So by that I mean \(\chi^{-1}(\chi(n))=\chi(\chi^{-1}(n))=n\) (I haven't really proven this to myself yet so I don't know if these are all true, just thought they were fun)

ganeshie8 (ganeshie8):

Very interesting! do we have : \[\sum\limits_{d\mid n} \mu(d) \chi(d) = \prod\limits_{\text{primes}} (1-\chi(p))\]

OpenStudy (kainui):

I'll denote the Dirichlet convolution of something with itself n times like this: \[f^{\star n}(a)\] So for example: \[f \star f \star f \star f = f^{\star 4}\] It looks like this is true: \[\mu^{\star n} (p^k) = (-1)^k\binom{n}{k}\] And since it's multiplicative, we have not too much trouble just finding the general form of the nth mobius convolution with itself.

OpenStudy (kainui):

@ganeshie8 Ahhh cool I think we might I'm still trying to work it out, I am also curious what \(\chi \star u\) will look like next as well. So far I am getting a weird result though haha \[(\mu \star \chi)(p^k) = \chi(p^k) ( 1 - \tfrac{1}{p} ) = \phi(\chi(p^k))\]

OpenStudy (kainui):

Oh I see I am calculating something slightly different than you, no wonder I am getting different results. I was doing this: \[(\mu \star \chi)(n) = \sum_{d|n} \mu(d) \chi ( \tfrac{n}{d})\] You were doing this: \[((\mu* \chi) \star u)(n) = \sum_{d|n} \mu(d) \chi (d)\]

ganeshie8 (ganeshie8):

yeah mine is not a straight convolution

ganeshie8 (ganeshie8):

More generally, for any multiplicative function \(f(n)\) we have: \[\sum\limits_{d\mid n} \mu(d)f(d) = \prod\limits_{\text{primes}} (1-f(d))\]

OpenStudy (kainui):

Awesome, I got it! Actually this just made me realize the same thing in a slightly different way that you've just written, really we can replace anywhere we find \(N(n)=n\) with \(\chi(n)\) and so it has all the same convolutions as it.

OpenStudy (kainui):

For fun I wanted to see what using the arithmetic derivative would give us for an upper bound under a certain peculiar circumstance... \[\chi(n) = \prod_{primes} p^{kp}\] \[\chi'(n) = \Omega(n)\chi(n)\] Now two new relations: \[\chi'(rad(n)) = \omega(n)\chi(rad(n))\] \[\chi'(n\#) = \pi(n) \chi(n \#)\] using the upper bound on the arithmetic derivative: \[\chi'(n\#) \le \chi(n \#)\frac{\log_2( \chi(n \#))}{2}\] \[\pi(n) \le \frac{\log_2( \chi(n \#))}{2}\] \[\pi(n) \le \frac{1}{2}\sum_{primes} p \log_2(p)\] However that's not really too exciting since we already know \[\pi(n) \le \sum_{primes} p\] Hahaha at least I tried. xD